# HiPreNets: High-Precision Neural Networks through Progressive Training

### v2

@@ 4 Loss Function Design for $L^{\infty}$ Norm Error Reduction
This choice of weighting term increases the importance of accurately fitting regions where the target function exhibits large magnitudes, which in the case of the refinement networks, corresponds to peaks in the target residuals. By emphasizing these high-magnitude regions, the refinement networks are encouraged to reduce errors where they matter most for controlling the $L^{\infty}$ norm.
We evaluate each function in Table 1 using a HiPreNet configuration with four stages and network widths [5–10–15–20], employing the tanh activation function and BFGS as the optimizer. For the refinement networks, we compare using the two different loss functions to train them: mean squared error (MSE) and weighted mean squared error (WMSE). The resulting test RMSE and $L^{\infty}$ norm errors are reported in Table 6.
@@
Table 6: Final RMSE and $L^{\infty}$ norm error for each function using MSE and WMSE loss functions in the HiPreNet framework.
Figure 8 shows the validation error from each network in the HiPreNet training process first when MSE is used as the loss function, and then when WMSE is used as the loss function. We observe that WMSE can achieve a lower $L^{\infty}$ norm error and even a lower RMSE than standard MSE loss, although the better choice between the two loss functions appears highly problem dependent.
For Function I.29.16 and I.13.12, using WMSE led to worse performance in both RMSE and $L^{\infty}$ norm error compared to standard MSE. This suggests that WMSE may overfit to regions that contain a small subset of data points with large weights. In such cases, the weighting scheme can misdirect learning capacity if residuals exhibit erratic or noisy behavior, causing the model to struggle to generalize. Moreover, as mentioned previously, the $L^{\infty}$ norm itself is inherently challenging to optimize due to its non-differentiability and sensitivity to outliers. WMSE serves as a proxy to mitigate this, but it cannot fully replicate direct $L^{\infty}$ norm minimization.
@@
Figure 8: Comparison of residual predictions at each stage for Function I.6.2 using MSE loss (left column) and WMSE loss (right column), both with a [5-10-15-20] configuration. Note that the axis and colorbar scales between neighboring plots may differ to accurately illustrate the loss landscape from each approach.
5 Weighted Sampling Based on Residual Errors
Even after the seemingly successful training of a neural network, certain regions of the input domain can remain poorly approximated, particularly near domain boundaries or in areas with sharp gradients. To address this, we introduce strategies that refine the training dataset based on a model’s error distribution. By prioritizing regions with large residuals, networks can focus their learning on where it is most needed, improving both global accuracy and worst-case error.
To further improve the performance of HiPreNet training, we introduce a weighted sampling strategy that leverages information from the current residuals to adaptively guide the training of subsequent networks. Rather than treating all training points equally, we assign greater sampling priority to data points with higher residual error, ensuring that regions with poor approximation are more strongly emphasized in future training. No new data is generated for this approach, it is a resampling of the existing training data.
@@ 5 Weighted Sampling Based on Residual Errors
Given a discrete set of training samples $\textbf{x}=\{x_{1},x_{2},\dots,x_{N}\}$, we define a sampling probability $p_{k}$ at each iteration $i$ for each training sample $_{k}\in\textbf{x}$ x where
@@
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$triangleright estimate of true $e_{0}$
@@
  $f^{NN}_{i}(\textbf{x})\leftarrow f_{i}^{NN}+e_{i-1}\hat{r}_{i-1}^{NN}(\textbf{x})$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@
To evaluate the method, we applied weighted sampling across all five benchmark functions from the Feynman dataset using four-stage HiPreNets with widths [5–10–15–20], averaging across three simulations with different random seeds. A weighted resampling of the dataset before training each refinement networks focuses them on handling more challenging regions of the dataset without increasing dataset size, offering a computationally efficient route to reducing error. The sampling strategy generally maintains good overall coverage of the domain while prioritizing regions with higher residual error. In particular, samples near the boundaries of the domain, where the residuals tend to be largest, are selected more frequently during training. This suggests that the weighted sampling method effectively concentrates training effort on the areas that are most difficult to approximate, without requiring an increase in dataset size.
@@
To summarize, this method can lead to a noticeable reduction in both RMSE and $L^{\infty}$ norm error. Similar to what was observed for the weighted mean square error loss function, weighted sampling does not guarantee a reduction in either error metric. It may cause the models to overfit to the training data, reducing their ability to generalize and worsening their performance on the validation data.
@@
Table 7: The final validation RMSE and $L^{\infty}$ norm error
for both MSE and WMSE loss functions using the weighted sampling technique.
Figure 9: Bar chart showing the best validation RMSE and $L^{\infty}$ norm error from Table 6 versus Table 7 for each function. To choose the best results, we choose the MSE or WMSE results based on which resulted in the lowest RMSE and lowest $L^{\infty}$ norm error for that given function. Table 6 uses the full dataset to train each refinement network, while Table 7 uses the weighted sampling technique to train each refinement network.
6 Targeted Model Correction with Local Residual Patching
In many real-world applications, models perform well on average but struggle with edge cases. For example, in medical diagnosis, a model may mispredict outcomes for patients with rare conditions, while in autonomous driving, unusual visual scenarios can lead to high prediction errors. For the functions we attempt to learn here, these edge cases are often along the boundary of the input domain. To address these challenges, we introduce a local patching technique that selectively refines model predictions in regions where performance is weakest.
In this approach, once the HiPreNet algorithm is complete after $m$ iterations, we evaluate the ensemble’s performance on a randomly selected set of validation points. We then identify the validation point $x_{max}$ with the largest absolute prediction error.
[…]
Around this point, we define a neighborhood $N(x_{max})$ by selecting all points $x$ within a specified radius $r$ such that
$N(x_{max})=\left\{x\in\textbf{x},||x-x_{max}||_{2}\leq r\right\}$
Choosing the value for this radius is currently done manually, which does introduce a challenge in balancing there being enough data to train the localized network on versus introducing overfitting to a localized region of the domain.
These nearby points have their associated residuals $r_{m}(N(x_{max}))$ normalized according to
$\hat{r}_{m}(x_{max})=\frac{r_{m}(N(x_{max}))}{\max|{r_{m}(N(x_{max}))}|}$
and form the training set for a new neural network $f_{patch}(x)$, specifically designed to model the behavior within this localized high-error region. After training, this network serves as a targeted “patching network” that refines the ensemble’s predictions by learning the normalized residuals in the identified area of weakness. The final prediction is now given by
$f^{NN}_{m+1}(\textbf{x})=f_{m}^{NN}(\textbf{x})+\mathbbm{1}(\textbf{x})f_{patch}(\textbf{x})$
$\mathbbm{1}(\textbf{x})=\begin{cases}1\quad\text{for }x\in N(x_{max})\\ 0\end{cases}$
As an example, we consider Function I.13.12. From the left plot in Figure 10, the maximum validation error after training a 5-10-15-20 residual ensemble is located at the peak in the lower right corner of the domain.
Figure 10: Validation data residuals for Function I.13.12: the left plot shows final residuals after standard HiPreNet training and the right plot includes the correction from a localized patching network.
The sample at this peak is chosen, as well as all samples in a radius of $r=0.75$ around it. A fully connected network with two hidden layers of 16 neurons each is then trained to learn the normalized residuals of these samples. After training this network, it is used to “patch” the region of the domain spanned by these samples. For any input to the final model, if it falls within this region, the patching network will be used and its prediction will be added to the final output. We then generate 1000000 new test data samples to test the generalization capabilities of this patching approach.
The RMSE and $L^{\infty}$ norm errors on the validation data and test data from before and after the patching network is applied are given in Table 8. The final validation data residuals after patching are shown in the right plot in Figure 10, and the final test data residuals before and after patching are given in Figure 11. As shown, by applying this patching approach, both the RMSE and the $L^{\infty}$ norm error can be reduced by nearly an order of magnitude.
Data
Initial
RMSE
Initial $L^{\infty}$
Norm Error
Patching
RMSE
Patching $L^{\infty}$
Norm Error
I.13.12 Validation 1.399e-07 1.116e-04 3.284e-08 2.270e-05
I.13.12 Test 2.396e-07 2.325e-04 3.934e-08 3.419e-05
@@ 6 Targeted Model Correction with Local Residual Patching
Table 8: Effect of applying local patching on Function I.13.12 with a comparison of validation RMSE and $L^{\infty}$ norm error before and after patching.
Figure 11: Test data residuals for Function I.13.12: the left plot shows final residuals after standard HiPreNet training and the right plot includes the correction from a localized patching network.
By targeting the specific area around the worst-performing validation point, the patching network provides a focused correction that improves both RMSE and the $L^{\infty}$ norm error. This technique demonstrates a scalable strategy for enhancing model reliability in applications where global accuracy masks localized weaknesses.

### v3

@@ 4.1 Loss Function Design for $L^{\infty}$ Norm Error Reduction
This choice of weighting term increases the importance of accurately fitting regions where the target function exhibits large magnitudes, which in the case of the refinement networks, corresponds to peaks in the target residuals. By emphasizing these high-magnitude regions, the refinement networks are encouraged to reduce errors where they matter most for controlling the $L^{\infty}$ norm. For our purposes, the offset prevents zero-weighting of small-residual points, and the linear form avoids overly concentrating on single outliers.
We evaluate each function in Table 1 using a HiPreNet configuration with four stages and network widths [5–10–15–20], employing the tanh activation function and BFGS as the optimizer. For the refinement networks, we compare using the two different loss functions to train them: mean squared error (MSE) and weighted mean squared error (WMSE). The resulting test RMSE and $L^{\infty}$ norm errors are reported in Table 6. We observe that WMSE can achieve a lower $L^{\infty}$ norm error and even a lower RMSE than standard MSE loss, although the better choice between the two loss functions appears highly problem dependent.
@@
Table 6: Final RMSE and $L^{\infty}$ norm error for each function using MSE and WMSE loss functions in the [5-10-15-20] HiPreNet configuration. Performance differences are highly problem-dependent.
In particular, for Function I.29.16 and I.13.12, using WMSE led to worse performance in both RMSE and $L^{\infty}$ norm error compared to standard MSE. This suggests that WMSE may overfit to regions that contain a small subset of data points with large weights. In such cases, the weighting scheme can misdirect learning capacity if residuals exhibit erratic or noisy behavior, causing the model to struggle to generalize. Moreover, as mentioned previously, the $L^{\infty}$ norm itself is inherently challenging to optimize due to its non-differentiability and sensitivity to outliers. WMSE serves as a proxy to mitigate this, but it cannot fully replicate direct $L^{\infty}$ norm minimization.
@@
4.2 Weighted Sampling Based on Residual Errors
To further improve the performance of HiPreNet training, we introduce a weighted sampling strategy that leverages information from the current residuals to adaptively guide the training of subsequent networks. Rather than treating all training points equally, we assign greater sampling priority to data points with higher residual error, ensuring that regions with poor approximation are more strongly emphasized in future training. No new data is generated for this approach, it is a resampling of the existing training data. This can also be thought of a data-driven way to choose the weight in the previously described WMSE loss function.
@@ 4.2 Weighted Sampling Based on Residual Errors
Given a discrete set of training samples $\textbf{x}=\{x_{1},x_{2},\dots,x_{N}\}$, we define a sampling probability $p_{k}$ at each iteration $i$ for each training sample $x_{k}\in\textbf{x}$ where
@@
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance toltol
@@
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$ $\triangleright$ estimate of true $e_{0}$
@@
  $f^{NN}_{i}(\textbf{x})\leftarrow f_{i-1}^{NN}+e_{i-1}\hat{r}_{i-1}^{NN}(\textbf{x})$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$ $\triangleright$ estimate of true $e_{i}$
@@
To evaluate the method, we applied weighted sampling across all five benchmark functions from the Feynman dataset using four-stage HiPreNets with widths [5–10–15–20], choosing the lowest error across three simulations with different random seeds. A weighted resampling of the dataset before training each refinement networks focuses them on handling more challenging regions of the dataset without increasing dataset size, offering a computationally efficient route to reducing error. The sampling strategy generally maintains good overall coverage of the domain while prioritizing regions with higher residual error. In particular, samples near the boundaries of the domain, where the residuals tend to be largest, are selected more frequently during training. This suggests that the weighted sampling method effectively concentrates training effort on the areas that are most difficult to approximate, without requiring an increase in dataset size.
@@
To summarize, this method can lead to a noticeable reduction in both RMSE and $L^{\infty}$ norm error. Similar to what was observed for the weighted mean square error loss function, weighted sampling does not guarantee a reduction in either error metric. It may cause the models to overfit to the training data, reducing their ability to generalize and worsening their performance on the validation data. Additionally, the weighted sampling technique is one instance of a broader principle: identify regions where the current model performs poorly and concentrate training resources there. In settings where new data can be generated cheaply, this principle can alternatively be implemented by augmenting the training dataset with new samples drawn from high error regions, rather than resampling existing data. We return to this idea in Section 6 in the context of a high dimensional power systems example, where the simulator allows cheap data generation and the high error region has a clear physical interpretation.
@@
Table 7: The final validation RMSE and $L^{\infty}$ norm error for both MSE and WMSE loss functions using the weighted sampling technique. Using weighted sampling with MSE loss results in lower error than using weighted sampling with WMSE loss, aside from for Function I.29.16.
Figure 8: Comparison of best-achieved validation RMSE and $L^{\infty}$ norm error across all five functions, with and without weighted sampling. For each function, the better of MSE and WMSE loss is shown. Weighted sampling reduces both error metrics for four of the five functions.
4.3 Targeted Model Correction with Local Residual Patching
For a given problem, even after the HiPreNet model has been fully trained, a dominant spike in the residuals may remain in a localized region of the input domain. This is apparent in Figure 9 (left), where the final residuals for a model trained to approximate Function I.13.12 exhibit a large peak on the boundary, corresponding to the location of the $L^{\infty}$ norm error. This observation motivates a targeted post-training correction, where rather than retraining the entire ensemble, we train a small “patching” network focused exclusively on the neighborhood of the spike.
In this approach, once the HiPreNet model is trained after $m$ iterations, we evaluate its performance on a validation dataset to “probe” the model’s weaknesses. With $r_{m}(x)=f(x)-f_{m}^{NN}(x)$ as the final residuals of the model, where $f(x)$ is the true target and $f_{m}^{NN}(x)$ is the model after $m$ iterations, we identify the validation point $x_{max}=\arg\max|r_{m}(x)|$ that has the maximum absolute residual. We next define a Gaussian weight function centered on $x_{max}$
$w(x)=\exp\left(-\frac{\|x-x_{max}\|_{2}^{2}}{2\sigma^{2}}\right)=\exp\left(-\frac{d^{2}}{2\sigma^{2}}\right)$
where $d=\|x-x_{max}\|_{2}$ is the distance from a given point $x$ to $x_{max}$ and $\sigma$ controls the extent of the neighborhood around the spike. We then look at only the neighborhood $N(x_{max})=\{x\in\text{validation dataset}\ |\ w(x)>\tau\}$ of the $K=|N(x_{max})|$ validation points around $x_{max}$ that satisfy $w(x)>\tau$ for a set threshold $\tau$. With this weight function, a patching network $f_{patch}(x)$ is trained on this neighborhood of points by using the loss function
$L_{patch}=\frac{1}{K}\sum_{i=1}^{K}w(x_{i})\left(\hat{r}_{m}(x_{i})-f_{patch}(x_{i})\right)^{2},\quad x_{i}\in N(x_{max})$
where $\hat{r}_{m}(x_{i})=r_{m}(x_{i})/\max(|r_{m}|)$ are the normalized final residuals. Once this network is trained, the final corrected prediction is
$f_{m}^{patch}(x)=f_{m}^{NN}(x)+\mathds{1}(x)\max(|r_{m}|)f_{patch}(x)$
where the indicator function
$\mathds{1}(x)=\begin{cases}1\ \text{if}\ x\in N(x_{max})\\ 0\ \text{otherwise}\end{cases}$
ensures that the patching network is applied only to inputs within the patching neighborhood $N(x_{max})$, leaving all other inputs unchanged.
The patching network $f_{patch}$ consists of two hidden layers of 16 neurons each with tanh activation functions, giving a total of 337 trainable parameters. The network is optimized using the BFGS algorithm with a gradient tolerance of $10^{-12}$ and a maximum of 20,000 iterations, consistent with the optimizer used for the HiPreNet refinement networks. Importantly, this network also uses a tanh activation applied to the output, constraining its predictions to [1,1][-1,1] for any input. This is a critical design choice with two important consequences. First, since the normalized training targets lie in [1,1][-1,1] by construction, the tanh does not restrict the network’s expressivity within its training distribution. Second, it provides a hard guarantee that the correction magnitude never exceeds $\max|r_{m}|$ for any input, including unseen test points. Without this constraint, we observed the network extrapolating to corrections orders of magnitude larger than intended on slightly different test inputs, catastrophically worsening predictions. We proceed to use this setup to test all of the following examples.
For this problem, we use $\sigma=0.5$ and $\tau=10^{-3}$, retaining approximately 21,000 of 1,000,000 validation points for training. The threshold $\tau=10^{-3}$ was chosen as a practical cutoff for computational efficiency. For any choice of $\sigma$, a point at distance $d$ from $x_{\max}$ is retained if
$\exp(-d^{2}/2\sigma^{2})>\tau,$
which corresponds to
$d\leq\sigma\sqrt{-2\ln\tau}.$
This means setting $\tau=10^{-3}$ retains all points within approximately $3.7\sigma$ of $x_{\max}$. As such, points with $w(x)<\tau$ contribute less than 0.1% of the maximum weight to the loss function, and their inclusion would increase training cost without meaningfully affecting the learned correction.
The parameter $\sigma$ was selected by sweeping over $\{0.10,0.25,0.50,0.75,1.00,1.50\}$ and examining the sensitivity of the results, shown in Table 8. Values of $\sigma$ below 0.25 retained too few points for reliable training and showed no improvement in validation $L^{\infty}$ norm error. The value of $\sigma=0.75$ was chosen as a representative value as it is the smallest value that retains enough points $n$ for reliable training and attains the lowest observed validation $L^{\infty}$ norm error.
$\sigma$ $n$ Val RMSE Val $L^{\infty}$ Norm Error
0.10 378 1.531e-08 1.356e-05
0.25 3,634 1.460e-08 1.384e-05
0.50 21,410 4.102e-09 1.726e-06
0.75 64,768 4.824e-09 1.559e-06
1.00 164,519 3.541e-09 1.559e-06
1.50 906,250 3.786e-09 1.559e-06
Table 8: Sensitivity of patching results to the parameter $\sigma$, showing validation RMSE and $L^{\infty}$ norm error alongside the number of training points retained above threshold $\tau=10^{-3}$. Based on these validation results, $\sigma=0.75$ was selected as it is the smallest value in the stable range that retains enough points for reliable training. Values of $\sigma$ below 0.25 retain too few points for reliable training.
We evaluate the selected configuration on an independently generated test dataset of 1,000,000 samples drawn from the same distribution as the validation data but with a different random seed, ensuring no overlap with the data used to train the patching network or select $\sigma$. The results are shown in Table 9 and illustrated in Figures 9 and 10. On the validation data, patching reduces RMSE from 2.758e-08 to 4.824e-09 and $L^{\infty}$ norm error from 2.293e-05 to 1.559e-06, which are approximately 83% and 93% decreases, respectively. This is visible in Figure 9 as a dramatic reduction in the residual spike. On the independently generated test data, the $L^{\infty}$ norm error is reduced from 2.127e-04 to 1.898e-04, a reduction of approximately 11%, with RMSE slightly decreasing from 2.271e-07 to 2.001e-07.
Data Model RMSE $L^{\infty}$ Norm Error
Validation HiPreNet Stage 4 2.758e-08 2.293e-05
Original Patching 4.824e-09 1.559e-06
Test HiPreNet Stage 4 2.271e-07 2.127e-04
Original Patching 2.001e-07 1.898e-04
@@ 4.3 Targeted Model Correction with Local Residual Patching
Table 9: RMSE and $L^{\infty}$ norm error for Function I.13.12 before and after local residual patching, evaluated on validation and test data.
Figure 9: Validation data residuals for Function I.13.12: the left plot shows final residuals after standard HiPreNet training and the right plot shows residuals after applying the localized patch network. The spike near the domain boundary, corresponding to the $L^{\infty}$ norm error, is reduced by nearly an order of magnitude Original Patching. Note that axis and colorbar scales differ across plots. The spike in the right plot is the location of the new $L^{\infty}$ norm error.
Figure 10: Test data residuals for Function I.13.12: the left plot shows final residuals after standard HiPreNet training and the right plot shows residuals after applying the localized patch network trained on validation data. The spike is partially reduced, with the $L^{\infty}$ norm error decreasing by approximately 11%. The smaller improvement relative to validation reflects the test spike being approximately $9\times$ larger in magnitude than the validation spike, which limits the correction achievable within the bounded output of the patching network. Note that axis and colorbar scales differ across plots.
The smaller improvement on test data reflects an inherent limitation of the technique. The $L^{\infty}$ norm error point on the test data was found to lie within the patch region, with a distance of $0.27$ from the $L^{\infty}$ norm error point on the validation data and a Gaussian weight of $w(x_{\max}^{test})=\exp(-0.27^{2}/(2\times 0.75^{2}))\approx 0.93>\tau=10^{-3}$. Even though the test spike is inside the patching region, applying the tanh activation to the network’s outputs guarantees that the correction magnitude never exceeds the magnitude of the validation spike, which is $\max|r_{m}|=2.293\text{e-05}$. The test spike is approximately $9\times$ larger at 2.127e-04, so the patch can only partially correct it. This indicates that the validation dataset is not large or robust enough to properly capture the spike, which suggests a method for targeted sampling if further data generation is allowed, which we investigate in subsection 4.3.2.
Returning to the assumption that no more data can be generated for training, we apply the patching technique to the HiPreNet models trained for each of the Feynman functions, and report the results in Table 11. For each function aside from Function I.29.16, we choose the $\sigma$ that retains the largest amount of samples under 100000 samples to ensure sufficient coverage of the patching region without introducing overfitting. For Function I.29.16, we choose $\sigma=0.75$ to demonstrate a potential failure case for the method. After applying patching, there is a drop in the test $L^{\infty}$ norm error for all functions aside from Function I.29.16, where it moderately increases. Of note, the test $L^{\infty}$ norm error for function I.6.2 is reduced by approximately 59%, and the test $L^{\infty}$ norm error for function I.26.2 is reduced by approximately 52%.
To understand why patching worsens performance on the test data for Function I.29.16, we examine the stability of the $L^{\infty}$ norm error location across independently generated validation datasets. Figures 11 and 12 show the $L^{\infty}$ norm error locations across five validation datasets in the normalized input space for Function I.13.12 and I.29.16, respectively. Unlike Function I.13.12, where the $L^{\infty}$ norm error point was consistently located near the same boundary region, Function I.29.16 exhibits two distinct clusters of $L^{\infty}$ norm error locations separated by a distance of approximately 5.3 in the normalized input space. The patch network is trained on whichever cluster appears in the validation data, but has no effect if the test worst error falls in the other cluster. As such, when the worst error location is not stable, the technique cannot reliably generalize to unseen test data using only one patching network, suggesting that an iterative training process to construct multiple patching networks is necessary.
4.3.1 Iterative Patching
To test this idea, we apply the iterative patching procedure to Function I.29.16. Beginning from the HiPreNet Stage 4 predictions, we identify the largest residual sample on the validation data and train a patching network as described above. We then update the running predictions on both validation and test data, and repeat the process, meaning at each subsequent iteration, the largest residual sample is re-identified from the updated residuals, and a new patching network is trained and applied. This continues for a fixed number of iterations $P$, resulting in a stack of patching networks $\{f_{patch}^{(p)}\}_{p=1}^{P}$ with associated scales $\{\max|r^{(p)}|\}$ and largest residual samples $\{x_{max}^{(p)}\}$. Then, the corrected prediction at iteration $p$ is
$f^{(p)}(x)=f^{(p-1)}(x)+\mathds{1}^{(p)}(x)\max\left|r^{(p-1)}\right|f_{patch}^{(p)}(x),$
where $r^{(p-1)}(x)=f(x)-f^{(p-1)}(x)$ is the residual after p1p-1 patches, $f^{(0)}=f_{m}^{NN}$ is the Stage 4 HiPreNet output, and $\mathds{1}^{(p)}$ is the indicator function for the $p$-th patch neighborhood $N(x_{max}^{(p)})$.
By construction, each patch network targets the current dominant spike in the residuals, and successive iterations address distinct high error regions as earlier ones are corrected. This is the behavior needed for Function I.29.16, where the $L^{\infty}$ norm error location is unstable across two different regions. If the first patch corrects the spike in one region, the second iteration’s largest error search will naturally migrate to the other region, and a second patch network can be trained and applied there.
The results of applying $P=2$ iterative patches to Function I.29.16 are shown in Table 10. On the validation data, the $L^{\infty}$ norm error decreases monotonically across iterations, with the largest reduction occurring in the first two patches as the two dominant error clusters are successively corrected. On the test data, which was not used at any stage of the patching procedure, the $L^{\infty}$ norm error also decreases across iterations, confirming that the technique generalizes beyond the validation set when the iterative procedure is used. The improvement on test data is more modest than on validation, consistent with the bounded-output guarantee limiting the magnitude of each correction to at most $\max|r^{(p-1)}|$ on the validation residuals, which may underestimate the true spike magnitude on the test data.
Data Model RMSE $L^{\infty}$ Norm Error
Validation Initial 6.4418e-04 3.1251e-02
Stage 2 8.6174e-05 8.9618e-03
Stage 3 3.4741e-05 3.4392e-03
Stage 4 1.2301e-05 1.2590e-03
Patch 1 1.2107e-05 9.4403e-04
Patch 2 1.1742e-05 9.3998e-04
Test Initial 6.4112e-04 3.1083e-02
Stage 2 8.5425e-05 9.0803e-03
Stage 3 3.4677e-05 2.8813e-03
Stage 4 1.2255e-05 9.5549e-04
Original Patching 1.1777e-05 8.9009e-04
Table 10: RMSE and $L^{\infty}$ norm error for Function I.29.16 across HiPreNet ensemble stages and iterative Gaussian patching, evaluated on validation and test data. Patch iterations are applied sequentially to the Stage 4 predictions; intermediate test results are not reported as the patch networks are selected using validation data only.
We note that the iterative procedure introduces no additional hyperparameters beyond the number of iterations $P$, a decay rate $\gamma\in(0,1]$, and the shared $\tau$ used in each step. At iteration $p$, the neighborhood width is set to $\sigma^{(p)}=\sigma_{0}\cdot\gamma^{p}$, where $\sigma_{0}$ is the initial value selected by the sweep described above. This decay ensures that successive patch networks are trained on progressively smaller neighborhoods, concentrating each correction more tightly around the newly identified worst-error point as the dominant spikes are resolved. The total computational cost scales linearly with $P$, since each iteration trains one additional patch network of fixed size using BFGS.
Function $\sigma$ Data Model RMSE $L^{\infty}$ Norm Error
I.6.2 (2D) 0.25 Validation HiPreNet Stage 4 1.678e-11 1.342e-09
Original Patching 1.502e-11 2.930e-10
Test HiPreNet Stage 4 1.698e-11 1.551e-09
Original Patching 1.507e-11 6.367e-10
I.9.18 (6D) 0.75 Validation HiPreNet Stage 4 1.363e-07 4.388e-05
Original Patching 1.175e-07 3.833e-05
Test HiPreNet Stage 4 2.364e-07 1.831e-04
Original Patching 2.202e-07 1.393e-04
I.13.12 (2D) 0.75 Validation HiPreNet Stage 4 2.758e-08 2.293e-05
Original Patching 4.824e-09 1.559e-06
Test HiPreNet Stage 4 2.271e-07 2.127e-04
Original Patching 2.001e-07 1.898e-04
I.26.2 (2D) 0.25 Validation HiPreNet Stage 4 2.431e-05 8.549e-03
Original Patching 1.357e-05 3.578e-03
Test HiPreNet Stage 4 2.977e-05 1.495e-02
Original Patching 1.600e-05 7.167e-03
I.29.16 (3D) 0.75 Validation HiPreNet Stage 4 1.230e-05 1.259e-03
Original Patching 1.209e-05 9.440e-04
Test HiPreNet Stage 4 1.226e-05 9.555e-04
Original Patching 1.217e-05 1.398e-03
Table 11: RMSE and $L^{\infty}$ norm error before and after local residual patching for all five Feynman functions, evaluated on validation and test data. Function dimensionality is left-aligned below the function name.
Figure 11: $L^{\infty}$ norm error locations across five independently generated validation datasets for Function I.13.12, plotted in the normalized input space. The locations cluster tightly near the same boundary region across all five datasets, indicating that the $L^{\infty}$ norm error location is stable and the patch network trained on one dataset will be applied in the correct region for unseen data.
Figure 12: $L^{\infty}$ norm error locations across five independently generated validation datasets for Function I.29.16, plotted in the $x_{1}-x_{2}$ plane of the normalized input space. The locations fall into two distinct clusters separated by a distance of approximately 5.3, corresponding to two different boundary corners of the domain. When the validation $L^{\infty}$ norm error falls in one cluster and the test $L^{\infty}$ norm error falls in the other, the patch network is applied in the wrong region and cannot improve performance.
4.3.2 Targeted Sampling
As mentioned previously, the limited gains in accuracy on the test dataset observed with Function I.13.12 appear to be a result of the validation dataset not being enough to completely capture the spike in the residuals. In other words, the small area selected for patching is generally noisier and more oscillatory than the whole large scale original dataset. This suggests that there needs to be a dense dataset in the patching area to allow the model to completely capture the spike. With the ability to generate more data, this prompts us to generate a new dataset comprised of data that falls within a selected patching area, and then train the patching network on this dataset.
In practice, we calculate $x_{max}$ and the neighborhood around it, $N(x_{max})$, as previously discussed. We then generate a dataset of 50000 samples that fall within the domain encapsulated by $N(x_{max})$ and train the patching network on this dataset. We compare the performance of this approach to the previous approach in Table 12 on the same datasets, where we observe that using this targeted sampling approach consistently leads to lower test $L^{\infty}$ norm error, aside from Function I.29.16, whose degradation in performance we discussed previously.
Function $\sigma$ Data Model RMSE $L^{\infty}$ Norm Error
I.6.2 (2D) 0.25 Validation Original Patching 1.502e-11 2.930e-10
Targeted Patching 1.529e-11 2.930e-10
Test Original Patching 1.507e-11 6.367e-10
Targeted Patching 1.526e-11 2.946e-10
I.9.18 (6D) 0.75 Validation Original Patching 1.175e-07 3.833e-05
Targeted Patching 1.206e-07 3.833e-05
Test Original Patching 2.202e-07 1.393e-04
Targeted Patching 1.793e-07 9.822e-05
I.13.12 (2D) 0.75 Validation Original Patching 4.824e-09 1.559e-06
Targeted Patching 1.346e-08 8.670e-06
Test Original Patching 2.001e-07 1.898e-04
Targeted Patching 1.110e-07 8.867e-05
I.26.2 (2D) 0.25 Validation Original Patching 1.357e-05 3.578e-03
Targeted Patching 1.325e-05 3.578e-03
Test Original Patching 1.600e-05 7.167e-03
Targeted Patching 1.508e-05 6.261e-03
I.29.16 (3D) 0.75 Validation Original Patching 1.209e-05 9.440e-04
Targeted Patching 1.209e-05 9.440e-04
Test Original Patching 1.217e-05 1.398e-03
Targeted Patching 1.217e-05 1.398e-03
Table 12: RMSE and $L^{\infty}$ norm error comparing standard and augmented local residual patching for all five Feynman functions, evaluated on validation and test data. Function dimensionality is left-aligned below the function name.